OpenStudy (sadaf):

Which equation is y = –6x2 + 3x + 2 rewritten in vertex form?

OpenStudy (loser66):

What do you mean by vertex form?

OpenStudy (mathmale):

The equation of a quadratic in vertex form looks like this : y-k=a(x-h)^2, where (h,k) is the vertex and a is a constant coefficient. To find the "vertex form" of your given equation, you'll need to re-write the given equation by "completing the square" of –6x2 + 3x. Given y = –6x^2 + 3x + 2, first factor –6x2 + 3x into the form -6(x+(3/6), or -6(x+(1/2). Complete the square. Questions? Please ask.